OpenStudy (anonymous):

Find the equation of the line through (2,8) and parallel to y=3x-3. Write the equation using function notation

OpenStudy (agent47):

parallel means they have the same slope, which is 3. y-8=3(x-2) y-8=3x-6 y=3x-6+8=3x+2 Function notation: f(x)=3x+2